# Performance and Non-adversarial Robustness of the Segment Anything Model 2 in Surgical Video Segmentation

### v1

@@ 1 Introduction
Accurate segmentation of surgical tools and anatomical structures in endoscopic videos is fundamental for computer-assisted interventions and surgical robotics applications [1], such as instrument tracking, surgical scene understanding, and automated surgical skill assessment [2]. However, real-world clinical scenarios present challenges for robust segmentation, including visual corruptions like smoke, bleeding, specular reflections, and low illumination that can substantially alter the appearance of surgical scenes [3, 4]. These challenging conditions, which are commonly encountered during actual procedures, can degrade segmentation performance and potentially compromise patient safety if not properly addressed [5]. While fully supervised deep learning (DL) models have achieved promising results on standard surgical video datasets under controlled conditions, they often struggle to generalize effectively to the diverse and dynamic environments of live surgeries [6], which highlights the need for more robust and adaptable segmentation approaches that can maintain high accuracy across a wide range of real-world surgical scenarios Recently, foundation models for image segmentation, particularly the Segment Anything Model (SAM), have demonstrated zero-shot capabilities when provided with appropriate user prompts [7]. SAM introduced a paradigm shift towards interactive segmentation, where a single model can be flexibly applied to diverse segmentation tasks through prompting. However, direct application of SAM to medical imaging domains like surgical videos has revealed performance gaps compared to fully supervised DL models trained specifically on medical data [8, 9]. Building upon SAM’s success, the recently proposed SAM-2 extends these zero-shot interactive segmentation capabilities to video data [10].
@@
In this paper, we present the first experimental study comprehensively evaluating SAM-2’s performance on surgical video data. We leverage the SegSTRONG-C dataset [11], which provides mock endoscopic sequences both with and without simulated visual corruptions like smoke, bleeding, and low brightness to assess SAM-2’s effectiveness on standard surgical videos as well as its robustness to challenging real-world conditions. Our study compares SAM-2 against the original SAM, medical variants of SAM that process video as per-frame image segmentation, and state-of-the-art fully supervised DL models. We explore various prompting strategies to investigate how SAM-2 can best leverage its temporal modeling capabilities. The major contributions are three-fold. Firstly, we conduct the first comprehensive evaluation of SAM-2 on surgical video data, including uncorrupted sequences and those with simulated visual corruptions. Secondly, we provide an in-depth comparison of SAM-2 against other SAM variants and fully supervised models, highlighting the benefits of SAM-2’s temporal modeling for surgical videos. Finally, we explore various prompting strategies to optimize SAM-2’s performance on surgical data, including frame-wise and frame-sparse prompting.
@@ 2 Methods
Frame-wise ($N=1$) SAM [7] 0.6897 0.7419 0.5643 0.6260 0.2281 0.8496 0.8791 0.7681 0.8134 0.4088
@@
Frame-sparse ($N=300$) SAM [7] 0.4889 0.5327 0.4180 0.4580 0.2321 0.6026 0.6132 0.5458 0.5746 0.3697
@@ 3 Experiments
SAM-2 demonstrates competitive or superior performance compared to fully-supervised DL models across most conditions. On regular surgical videos, SAM-2 demonstrates competitive performance compared to them. With frame-wise prompting (i.e., $N=1$), SAM-2 achieves a DSC of $0.8543$ in a zero-shot manner, which is comparable to the best-performing fully-supervised model, UNet with DSC of $0.9372$. Notably, when using frame-sparse prompting ($N=300$), SAM-2’s performance improves, reaching a DSC of $0.9096$, which is very close to UNet’s performance and can outperform DeepLabv3+ and SegFormer.
@@
Under the frame-sparse setting ($N=300$), SAM-2 consistently outperforms the original SAM and its medical variants across all five conditions. The performance gap is particularly notable in challenging three conditions. For instance, in the bleeding scenario, SAM-2 achieves a DSC of $0.8628$, surpassing SAM ($0.5458$), MedSAM ($0.1276$), and SAM-Med2D ($0.4400$). Similarly, in smoky conditions, SAM-2 maintains high DSC performance ($0.8660$) compared to SAM ($0.5746$), MedSAM ($0.1102$), and SAM-Med2D ($0.5746$). Even in regular conditions, SAM-2 outperforms its counterparts by a large margin. The original SAM performs reasonably well with frame-wise prompting but degrades significantly with frame-sparse prompting, highlighting SAM-2’s better temporal modeling capabilities.
@@
Frame-sparse prompting ($N=300$) consistently improves SAM-2’s performance compared to frame-wise prompting ($N=1$) across all conditions. This improvement is substantial, with DSC increases of 6.47%, 1.71%, 6.07%, 9.25%, and -5.26% for regular, background change, bleeding, smoke, and low brightness conditions, respectively. These results suggest that using fewer prompts can actually lead to better performance for SAM-2 on the surgical videos from SegSTRONG-C dataset [11]. This counter-intuitive finding can be attributed to the contribution of SAM-2’s memory encoder and bank in leveraging temporal information. In contrast, frame-wise prompting may not consistently select optimal prompts, potentially introducing noise or inconsistencies in the segmentation process. By allowing SAM-2 to rely more on its temporal modeling capabilities, frame-sparse prompting enables the model to maintain coherence across frames while reducing the impact of suboptimal individual prompts. Moreover, SAM [7] and SAM-Med2D [9] show decreased performance with frame-sparse prompting, underscoring the unique advantage of SAM-2’s architecture in handling temporal data.

### v2

@@ 1 Introduction
Accurate segmentation of surgical tools and anatomical structures in endoscopic videos is fundamental for computer-assisted interventions and surgical robotics applications [1], such as instrument tracking, surgical scene understanding, and automated surgical skill assessment [2]. However, real-world clinical scenarios present challenges for robust segmentation, including visual corruptions like smoke, bleeding, specular reflections, and low illumination that can substantially alter the appearance of surgical scenes [3, 4]. These challenging conditions, which are commonly encountered during actual procedures, can degrade segmentation performance and potentially compromise patient safety if not properly addressed [5]. While fully supervised deep learning (DL) models have achieved promising results on standard surgical video datasets under controlled conditions, they often struggle to generalize effectively to the diverse and dynamic environments of live surgeries [6], which highlights the need for more robust and adaptable segmentation approaches that can maintain high accuracy across a wide range of real-world surgical scenarios Recently, foundation models for image segmentation, particularly the Segment Anything Model (SAM), have demonstrated zero-shot capabilities when provided with appropriate user prompts [7]. SAM introduced a paradigm shift towards interactive segmentation, where a single model can be flexibly applied to diverse segmentation tasks through prompting. However, direct application of SAM to medical imaging domains like surgical videos has revealed performance gaps compared to fully supervised DL models trained specifically on medical data [8, 9]. Building upon SAM’s success, the recently proposed SAM-2 extends these zero-shot interactive segmentation capabilities to video data [10]. SAM-2 introduces architectural modifications to handle temporal information in video data, including a memory encoder and memory bank to maintain consistency across video frames. These enhancements make SAM-2 a promising candidate for robust surgical video segmentation, as it can potentially capture the dynamic nature of surgical procedures.
@@
In this paper, we present the first experimental study comprehensively evaluating SAM-2’s performance on surgical video data. We leverage the SegSTRONG-C dataset [11], which provides mock endoscopic sequences both with and without simulated visual corruptions like smoke, bleeding, and low brightness to assess SAM-2’s effectiveness on standard surgical videos as well as its robustness to challenging real-world conditions. Our study compares SAM-2 against the original SAM, medical variants of SAM that process video as per-frame image segmentation, and state-of-the-art fully supervised DL models. We explore various prompting strategies to investigate how SAM-2 can best leverage its temporal modeling capabilities. Our experiments reveal SAM-2’s strengths in maintaining temporal consistency across video frames and its adaptability to surgical scenarios, even outperforming fully supervised models in some cases. The major contributions are three-fold. Firstly, we conduct the first comprehensive evaluation of SAM-2 on surgical video data, including uncorrupted sequences and those with simulated visual corruptions. Secondly, we provide an in-depth comparison of SAM-2 against other SAM variants and fully supervised models, highlighting the benefits of SAM-2’s temporal modeling for surgical videos. Finally, we explore various prompting strategies to optimize SAM-2’s performance on surgical data, including frame-wise and frame-sparse prompting.
@@ 2 Methods
Frame-wise Point Prompt ($N=1$) SAM [7] 0.6897 0.7419 0.5643 0.6260 0.2281 0.8496 0.8791 0.7681 0.8134 0.4088
@@
Frame-wise BBox Prompt ($N=1$) SAM [7] 0.6228 0.6272 0.5846 0.5884 0.2586 0.7952 0.7956 0.7496 0.7475 0.4846
MedSAM [8] 0.2423 0.3672 0.2347 0.2304 0.1577 0.4604 0.6209 0.4743 0.4582 0.4762
SAM-Med2D [9] 0.3377 0.3398 0.2748 0.2891 0.2314 0.5055 0.4389 0.3448 0.4689 0.4143
Frame-sparse Point Prompt ($N=10$) SAM [7] 0.7396 0.6935 0.6852 0.7128 0.2890 0.8539 0.8307 0.8180 0.8394 0.5221
MedSAM [8] 0.1284 0.1270 0.1302 0.1185 0.1192 0.1779 0.2183 0.2108 0.1797 0.1806
SAM-Med2D [9] 0.2787 0.3092 0.2440 0.2624 0.2671 0.5203 0.6003 0.4503 0.5006 0.4627
SAM-2 [10] 0.8182 0.9023 0.7982 0.7548 0.2570 0.8935 0.9267 0.8714 0.8484 0.4335
Frame-sparse Bbox Prompt ($N=10$) SAM [7] 0.6108 0.6259 0.5307 0.5458 0.2554 0.7721 0.7766 0.7125 0.7230 0.4505
MedSAM [8] 0.1963 0.3372 0.1881 0.1973 0.1665 0.4519 0.5902 0.4606 0.4560 0.4630
SAM-Med2D [9] 0.2687 0.2686 0.2337 0.2316 0.2313 0.4473 0.3862 0.3004 0.4033 0.4062
Frame-sparse Point Prompt ($N=300$) SAM [7] 0.4889 0.5327 0.4180 0.4580 0.2321 0.6026 0.6132 0.5458 0.5746 0.3697
@@ 3 Experiments
SAM-2 demonstrates competitive or superior performance compared to fully-supervised DL models across most conditions. On regular surgical videos, SAM-2 demonstrates competitive performance compared to them. With frame-wise point prompting (i.e., $N=1$), SAM-2 achieves a DSC of $0.8543$ in a zero-shot manner, which is comparable to the best-performing fully-supervised model, UNet with DSC of $0.9372$. Notably, when using frame-sparse point prompting ($N=300$), SAM-2’s performance improves, reaching a DSC of $0.9096$, which is very close to UNet’s performance and can outperform DeepLabv3+ and SegFormer.
@@
Under the frame-sparse point prompt setting ($N=300$), SAM-2 consistently outperforms the original SAM and its medical variants across all five conditions. The performance gap is particularly notable in challenging three conditions. For instance, in the bleeding scenario, SAM-2 achieves a DSC of $0.8628$, surpassing SAM ($0.5458$), MedSAM ($0.1276$), and SAM-Med2D ($0.4400$). Similarly, in smoky conditions, SAM-2 maintains high DSC performance ($0.8660$) compared to SAM ($0.5746$), MedSAM ($0.1102$), and SAM-Med2D ($0.5746$). Even in regular conditions, SAM-2 outperforms its counterparts by a large margin. The original SAM performs reasonably well with frame-wise prompting but degrades significantly with frame-sparse prompting, highlighting SAM-2’s better temporal modeling capabilities.
@@
Frame-sparse point prompting ($N=300$) consistently improves SAM-2’s performance compared to frame-wise point prompting ($N=1$) across all conditions. This improvement is substantial, with DSC increases of 6.47%, 1.71%, 6.07%, 9.25%, and -5.26% for regular, background change, bleeding, smoke, and low brightness conditions, respectively. These results suggest that using fewer prompts can actually lead to better performance for SAM-2 on the surgical videos from SegSTRONG-C dataset [11]. This counter-intuitive finding can be attributed to the contribution of SAM-2’s memory encoder and bank in leveraging temporal information. In contrast, frame-wise point prompting may not consistently select optimal prompts, potentially introducing noise or inconsistencies in the segmentation process. By allowing SAM-2 to rely more on its temporal modeling capabilities, frame-sparse prompting enables the model to maintain coherence across frames while reducing the impact of suboptimal individual prompts. Moreover, SAM [7] and SAM-Med2D [9] show decreased performance with frame-sparse prompting, underscoring the unique advantage of SAM-2’s architecture in handling temporal data.